# Entanglement between dependent degrees of freedom: Quasi-particle correlations

### v1

@@ Abstract
Common notions of entanglement are based on well-separated subsystems. However, obtaining such independent degrees of freedom is not always possible because of physical constraints. In this work, we explore the notion of entanglement in the context of dependent degrees of freedom. As a physically relevant application, we specifically study quantum correlation features for quasi-particle descriptions. Those are paramount for interacting light-matter systems, utilizing excitations of fermion-boson hybrid modes. By comparing independent and dependent degrees of freedom, we uncover that certain states are non-entangled although they would be entangled when only focusing on the common, independent description, and vice versa. Therefore, new insight is provided into the resourcefulness of quantum correlations within the rarely discussed context of dependent degrees of freedom for light-matter links in quantum information applications.
@@ I Introduction
Entanglement in hybrid systems [11, 12] is crucial for quantum links between distinct physical platforms, e.g., using matter properties for processing and light for communication [13, 12]. The quasi-particle description is a convenient way to describe such hybrid systems, including the assessment of quantum resources stemming from light-matter interactions [15, 14]. Interfacing bosonic and fermionic parts in quantum links, quasi-particles thus constitute joint excitations of quantized bosonic and fermionic modes resulting in mixed quantum statistics for the quasi-particles [16].
@@
Quantum properties of quasi-particles, including entanglement, have been studied [17, 15, 18, 19, 20, 14]. Still, the versatile nature of entanglement means that different forms of entanglement exist [21, 22, 23, 24]. For example, it can occur independently for particles and fields to the point that both forms can be classified as two distinct phenomena [25]. In addition, entanglement can be independent of the decomposition of the full system into quantized basis modes, allowing entanglement to persist regardless the chosen mode basis [26]. This and other findings raise the question of whether the common analysis of entanglement applies to the quasi-particle pictures, too.
@@
In optics, for example, commonly studied degrees of freedom include, but are not limited to orbital angular momentum, time-bin modes, spatial modes, and polarization [27]. In all these instances, entanglement is formulated in terms of independent degrees of freedom of purely bosonic excitations [28, 29, 30, 31, 32]. However, degrees of freedom can be dependent because of physical constraints, which is not commonly discussed and the focus of this work.
@@
In this work, we overcome the aforementioned limitation by characterizing entangled and non-entangled states between dependent degrees of freedom. This is mainly done for light-matter interfaces and their quasi-particle descriptions via a mode transformation to quasi-particle excitations arising from diagonalizing a Hamiltonian in terms of quantized fields. We show instances of both entangled and non-entangled states for dependent degrees of freedom whose independent counterparts behave oppositely, showing the distinctiveness of quantum correlations in such scenarios. Also, we compare the fermion-boson quasi-particle modes with analogous models consisting purely of bosonic and fermionic subsystems. Finally, a generalization to other scenarios and instances of multipartite entanglement is outlined.
@@
The remainder of the work is structured as follows. In Sec. II, we introduce the specific system under study. Entanglement for quasi-particle descriptions is established and investigated in Sec. III. A comparison with independent degrees of freedom, i.e., purely bosonic and purely fermionic scenarios, is discussed in Sec. IV. In Sec. V, we formalize and generalize our approach. We conclude in Sec. VI.
II Light-matter interface
@@
II.1 Quantized eigenmodes of a Hamiltonian
Consider a composite fermion-boson system in second quantization with independent field operators $\hat{f}$ and $\hat{b}$, respectively, obeying $\{\hat{f},\hat{f}\}=0=[\hat{b},\hat{b}]$ and $\{\hat{f},\hat{f}^{\dagger}\}=\hat{1}=[\hat{b},\hat{b}^{\dagger}]$, where $\hat{1}$ denotes the identity. The system’s Hamiltonian under study consists of free-field contributions and couplings, analogous to the Jaynes–Cummings model [33], reading
@@ II.1 Quantized eigenmodes of a Hamiltonian
with $\Delta=\sqrt{\delta^{2}+|\kappa|^{2}}$ and $\varphi^{2}+\beta^{2}=1$. Excitations of these eigenmodes define quasi-particles, which we later discuss in detail. See also Ref. [34] in this context. Furthermore, we can recast the Hamiltonian in Eq. (1) in a noninteracting form as
@@
and without cross-terms between the quasi-particle modes; see also Fig. 1 for the energies pertaining to the eigenmodes.
Figure 1: Energy bands $E_{\pm}$ from diagonalization of the Hamiltonian via the introduction of quasi-particle modes $\hat{p}_{\pm}$ in Eq. (5) as a function of the ratio of coupling strength $|\kappa|$ and detuning $\delta$.
II.2 Quantum states of the system
Via the vacuum state $|\mathrm{vac}\rangle$, the Fock basis in the boson-fermion description consists of the elements
@@ II.2 Quantum states of the system
The notion of entanglement in the original $|m,n\rangle_{\mathrm{FB}}$ is given in terms of independent degrees of freedom. However, we are going to explore entanglement with respect to the eigenmodes of the Hamiltonian, in which we separate quasiparticles excited through $\hat{p}_{+}$ from the excitations of $\hat{p}_{-}$. This is going to lead us to the notion of entanglement between dependent degrees of freedom.
III Quasi-particle entanglement
In this section, we establish and analyze the notion of entanglement for dependent degrees of freedom in the context of the model under study. We show how linear dependence plays a central role in this concept. Also, we explore how the different notions of entanglement for states of a given total particle number behave differently for independent and dependent degrees of freedom.
III.1 Product states in fermion-boson hybrid descriptions
For defining entanglement of dependent degrees of freedom, we firstly need to recapitulate non-entangled states for independent degrees of freedom. Specifically, product states of independent fermions and bosons take the form
@@ III.1 Product states in fermion-boson hybrid descriptions
where $\tilde{x}_{m}=x_{m}$ and $\tilde{y}_{n}=y_{n}/\sqrt{n!}$ relate expansion coefficients in Fock basis and excitations of quantized fields. As examples, the basis states in Eq. (6) are classical states with respect to the fermion-boson separation. A state $|\psi\rangle$ that cannot be expressed as defined in Eq. (10) is, by definition, entangled.
Note that, throughout most parts of this work, we focus on pure states. However, it is worth mentioning that classically correlated, i.e., separable states [36], can be introduced by using an ensemble of pure product states in form of a density matrix. Still, all concepts discussed here can be extended to mixed, inseparable states, as discussed in Sec. V.
III.2 Field-product states for dependent degrees of freedom
Generalizing the second line in Eq. (10) renders it possible to define non-entangled states for the quasi-particle modes $\hat{p}_{\pm}$, Eq. (3). That is, a state is said to be non-entangled in this $\pm$ description if it can be expressed in a factorized manner as follows:
@@ III.2 Field-product states for dependent degrees of freedom
with complex coefficients $x_{m}$ and $y_{n}$ such that the state is normalized, ${}_{\pm}\langle x,y|x,y\rangle_{\pm}=1$. Again, a convex mixture in terms of a density matrix extends this notion to separable mixed states in the $\pm$ separation.
Similarly to the previous case, and according to the definition in Eq. (11), the following $N$-particle states, where $N=m+n$, are product states:
@@
Note that, unsurprisingly, such states are generally product states with respect to the quasi-particle modes $\hat{p}_{\pm}$ but entangled with respect to the fermion-boson separation $\mathrm{FB}$ since a global mode transformation is used in Eq. (3). Also, we can choose $m,n\in\mathbb{N}$ since neither $\hat{p}_{+}^{{\dagger}m}$ nor $\hat{p}_{-}^{{\dagger}n}$ become zero for a finite integer, contrasting pure fermions.
III.3 Linear dependence and disentangled NOON states
It is straightforward to observe that, for any two pairs of integers $(m,n)$ and $(m^{\prime},n^{\prime})$, which obey $(m\varphi^{2}-n\beta^{2})/\sqrt{m+n}\neq(m^{\prime}\varphi^{2}-n^{\prime}\beta^{2})/\sqrt{m^{\prime}+n^{\prime}}$, the states $|m,n\rangle_{\pm}$ and $|m^{\prime},n^{\prime}\rangle_{\pm}$ are linearly independent. However, for a fixed $N>0$, there are only two linearly independent elements in total since the $\mathrm{FB}$ expansion is based on two basis elements only; see Eq. (13). Therefore, the identities
@@ III.3 Linear dependence and disentangled NOON states
with $\psi_{0}\psi_{N}\neq 0$, are not necessarily entangled for $N>2$. Recall that, by contrast, all NOON states are entangled for independent degrees of freedom. Here, whenever the coefficients $\psi_{0}$ and $\psi_{N}$ take the form of Eq. (16), such NOON states clearly admit a product form $|m,n\rangle_{\pm}$ for the dependent degrees of freedom $\pm$. Again, we emphasize that this disentanglement of quasi-particle NOON states is distinct from common discussions of entanglement. See Fig. 2
Figure 2: NOON-type states from Eq. (17) on the circle in an FB coordinate system, for $N=5$ quasi-particles, where $\varphi=\beta=1/\sqrt{2}$ and $\vartheta=0$. All non-entangled states in the $\pm$ description, Eq. (16), are shown as bullet points, including a $\pm$ coordinate system of the dependent degrees of freedom, which is indicated through dashed lines. Note that $N+1=6$ states, plus their reflections about the origin, are shown in the real plane, ignoring complex phases of $\psi_{0}$ and $\psi_{N}$ for simplicity. For independent degrees of freedom, all but the states $|N,0\rangle_{\pm}$ and $|0,N\rangle_{\pm}$ (and their reflections $-|N,0\rangle_{\pm}$ and $-|0,N\rangle_{\pm}$) alone would be entangled.
III.4 Quantum-correlated eigenstates of non-interacting Hamiltonians
Next, consider the non-interacting case, where $\kappa=0$ for the Hamiltonian in Eq. (1), and the independent degrees of freedom $\mathrm{FB}$. This yields a Hamiltonian which decomposes into the single-mode parts $\hbar\omega_{F}\hat{f}^{\dagger}\hat{f}$ and $\hbar\omega_{B}\hat{b}^{\dagger}\hat{b}$, as well as vanishing interactions, $\hat{H}_{\mathrm{int}}=0$. In that case, the eigenstates are necessarily products, $|m,n\rangle_{\mathrm{FB}}$ [Eq. (6)], for the non-degenerate case $\delta\neq 0$. This is generally true for independent degrees of freedom and $\hat{H}=\hat{H}_{1}\otimes\hat{1}_{2}+\hat{1}_{1}\otimes\hat{H}_{2}$, with subsystem identities $\hat{1}_{j}$ and local Hamiltonians $\hat{H}_{j}$ for $j\in\{1,2\}$. Further, the quasi-particle modes $\hat{p}_{\pm}$ were introduced to achieve a non-interacting representation of the Hamiltonian for $\kappa\neq 0$, Eq. (5). So the expectation derived from independent degrees of freedom would be that the $N$-particle eigenstates in Eq. (7) are products with respect to the $\pm$ separation.

### v2

# Entanglement between dependent degrees of freedom: Quasiparticle correlations

@@ Abstract
Common notions of entanglement are based on well-separated subsystems. However, obtaining such independent degrees of freedom is not always possible because of physical constraints. In this work, we explore the notion of entanglement in the context of dependent degrees of freedom. As a physically relevant application, we specifically study quantum correlation features for quasiparticle descriptions. Those are paramount for interacting light-matter systems, utilizing excitations of fermion-boson hybrid modes. By comparing independent and dependent degrees of freedom, we uncover that certain states are nonentangled although they would be entangled when only focusing on the common, independent description, and vice versa. Therefore, insight is provided into the resourcefulness of quantum correlations within the rarely discussed context of dependent degrees of freedom for light-matter links in quantum information applications.
@@ I Introduction
Entanglement in hybrid systems [11, 12] is crucial for quantum links between distinct physical platforms, e.g., using matter properties for processing and light for communication [13, 12]. The quasiparticle description is a convenient way to describe such hybrid systems, including the assessment of quantum resources stemming from light-matter interactions [15, 14]. Interfacing bosonic and fermionic parts in quantum links, quasiparticles thus constitute joint excitations of quantized bosonic and fermionic modes resulting in mixed quantum statistics for the quasiparticles [16].
@@
Various quantum properties of quasiparticles, including entanglement, have been studied [17, 15, 18, 19, 20, 14]. Still, the versatile nature of entanglement means that a plethora of different forms of entanglement exist [21, 22, 23, 24]. For example, it can occur independently for particles and fields to the point that both forms can be classified as two distinct phenomena [25]. In addition, entanglement can be independent of the decomposition of the full system into quantized basis modes, allowing entanglement to persist regardless of the choice of a mode basis [26]. This and other findings raise the question of whether the common analysis of entanglement applies to the quasiparticle picture, too.
@@
In optics, for example, commonly studied degrees of freedom include, but are not limited to orbital angular momentum, time-bin modes, spatial modes, and polarization [27]. In all these instances, entanglement is formulated in terms of independent degrees of freedom of purely bosonic excitations [28, 29, 30, 31, 32]. However, degrees of freedom can be dependent because of physical constraints, which is not commonly discussed and, therefore, is the focus of this work.
@@
In this work, we overcome the aforementioned limitation by characterizing entangled and nonentangled states between dependent degrees of freedom. This is mainly done for light-matter interfaces and their quasiparticle descriptions via a mode transformation to quasiparticle excitations arising from diagonalizing a Hamiltonian in terms of quantized fields. We show instances of both entangled and nonentangled states for dependent degrees of freedom whose independent counterparts behave oppositely, showing the distinctiveness of quantum correlations in such scenarios. Also, we compare the fermion-boson quasiparticle modes with analogous models consisting purely of bosonic and fermionic subsystems. Finally, a generalization to other scenarios and instances of multipartite entanglement is outlined.
@@
The remainder of the work is structured as follows. The concepts of dependent and independent degrees of freedom in the context of quantum systems are established in Sec. II. In Sec. III, we introduce the specific light-matter system under study. Entanglement for quasiparticle descriptions is established and investigated in Sec. IV. A comparison with independent degrees of freedom, specifically, purely bosonic and purely fermionic scenarios, is discussed in Sec. V. In Sec. VI, we discuss generalizations of our approach that apply to an arbitrary number of bosonic and fermionic modes. We conclude in Sec. VII.
II Independent and dependent degrees of freedom for composite quantum systems
The classification of quantum-physical degrees of freedom is motivated by the notion of statistical independence in probability theory. In this classical context, the ability to factorize a joint distribution of random variables into distributions for each variable is at the core of this concept. Based on the notion of statistical independence, a classification of dependent and independent degrees of freedoms can be introduced within quantum physics.
For a composite quantum system, consisting of Hilbert spaces $\mathcal{H}_{X}$ and $\mathcal{H}_{Y}$, the principles of quantum physics for independent degrees of freedom require the joint system to read $\mathcal{H}_{X}\otimes\mathcal{H}_{Y}$. This space includes factorizable states $|x\rangle_{X}\otimes|y\rangle_{Y}$, as well as all linear combinations thereof. If this construction applies, we speak about independent degrees of freedom in which the common notion of entanglement is well defined; otherwise, we refer to $X$ and $Y$ as dependent.
In the case of physical constraints limiting the available state space, one has to treat the scenario more carefully. This applies particularly if joint restrictions apply across the $X$ and $Y$ parts of the composite systems. For instance, the symmetrization postulate generally does not admit product states $|x\rangle_{X}\otimes|y\rangle_{Y}$ for identical particles $X$ and $Y$ in first quantization. Rather, the (anti)symmetrized states $|x\rangle_{X}\otimes|y\rangle_{Y}\pm|y\rangle_{X}\otimes|x\rangle_{Y}\pm$ form the basis of the state space for the now dependent degrees of freedom. Specifically note that symmetric and antisymmetric states are obtained via projections onto the symmetric and antisymmetric subspaces of the composite quantum system. This impacts the very definition of entanglement (see, e.g., Ref. [33] and the references therein).
As we are going to explore in this work via a detailed study of quasiparticle excitations in second quantization, the entanglement characteristics of quasiparticles describe another example of dependent degrees of freedom. For instance, we show that nonentangled states are projected product states [see Eq. (35)]. Based on this result and the example mentioned above, we formalize this approach. Let $\mathcal{H}=\hat{P}(\mathcal{H}_{X}\otimes\mathcal{H}_{Y})$ be a composite Hilbert space subject to a constraint given by the projection $\hat{P}$, with $\hat{P}^{2}=\hat{P}=\hat{P}^{\dagger}$. A pure nonentangled state can be defined as
where $|x\rangle_{X}\in\mathcal{H}_{X}$ and $|y\rangle_{Y}\in\mathcal{H}_{Y}$ are local states of the joint system. Note that this definition is identical with the one for a factorizable state for independent degrees of freedom, where $\hat{P}=\hat{1}$. An entangled state $|\Psi\rangle$ does not admit the decomposition from Eq. (1), $|\Psi\rangle\neq\hat{P}\,(|x\rangle_{X}\otimes|y\rangle_{Y})$. Of course, we choose $\langle\Psi|\Psi\rangle=1$ such that the state is normalized after projection.
Furthermore, a mixed, only classically correlated state takes the form
[…]
and a state that does not obey this form is consequently inseparable, i.e., entangled. Also note that this definition straightforwardly generalizes to multipartite systems, and it is compatible with the well-established notion of separability and inseparability for independent degrees of freedom [34] when $\hat{P}=\hat{1}$ holds true. Therefore, a separable state is here defined as a state that obeys the physical constraint (given by the projector $\hat{P}$) and only contains classical correlations (via the joint probability distribution $P$) of tensor-product states $|x\rangle_{X}\otimes|y\rangle_{Y}$.
In addition to the quasiparticle description we explore in the remainder of this work, we draw a comparison to fundamental observables. Consider the degree of freedom pertaining to the total angular momentum, $J^{2}$, and one for the angular momentum in the $z$ direction, $J_{z}$. It is well known that the corresponding angular-momentum operators can be jointly diagonalized,
$\text{with}\quad-j\leq m\leq j$ $\quad\text{and}\quad j\geq 0.$
where $j$ and $m$ could, in principle, take arbitrary (half) integer values without the restrictions. The projection $\hat{P}$ then makes sure that the angular momentum in one direction never exceeds the total angular momentum, $|m|\ngtr j$. The above definition further allows us to define nonentangled states as follows:
For the remainder of this work, we compare entanglement for dependent and independent degrees of freedom in the context of light-matter interfaces. Those systems are of particular interest from the physical as well as from the quantum-informational perspective, as outlined in the Introduction.
III Light-matter interface
@@
III.1 Quantized eigenmodes of a Hamiltonian
Consider a composite fermion-boson system in second quantization with independent field operators $\hat{f}$ and $\hat{b}$, respectively, obeying $\{\hat{f},\hat{f}\}=0=[\hat{b},\hat{b}]$ and $\{\hat{f},\hat{f}^{\dagger}\}=\hat{1}=[\hat{b},\hat{b}^{\dagger}]$, where $\hat{1}$ denotes the identity. The system’s Hamiltonian under study consists of free-field contributions and couplings, analogous to the Jaynes–Cummings model [36], reading as
@@ III.1 Quantized eigenmodes of a Hamiltonian
with $\Delta=\sqrt{\delta^{2}+|\kappa|^{2}}$ and $\varphi^{2}+\beta^{2}=1$. Excitations of these eigenmodes define quasiparticles, which we later discuss in detail. See also Ref. [37] in this context. Furthermore, we can recast the Hamiltonian in Eq. (6) in a noninteracting form as
@@
and without cross-terms between the quasiparticle modes; see also Fig. 1 for the energies pertaining to the eigenmodes.
Figure 1: Energy bands $E_{\pm}$ from diagonalization of the Hamiltonian via the introduction of quasiparticle modes $\hat{p}_{\pm}$ in Eq. (10) as a function of the ratio of the coupling strength $|\kappa|$ and the detuning $\delta$.
III.2 Quantum states of the system
Beginning with the vacuum state $|\mathrm{vac}\rangle$, the Fock basis in the boson-fermion description consists of the elements
@@ III.2 Quantum states of the system
The notion of entanglement in the original $|m,n\rangle_{\mathrm{FB}}$ is given in terms of independent degrees of freedom. However, we are going to explore entanglement with respect to the eigenmodes of the Hamiltonian, in which we separate quasiparticles excited through $\hat{p}_{+}$ from the excitations of $\hat{p}_{-}$, resulting in dependent degrees of freedom.
IV Quasiparticle entanglement
In this section, we analyze the notion of entanglement for dependent degrees of freedom for the model under study. We show how linear dependence plays a central role in this concept. Also, we explore how the different notions of entanglement for states of a given total particle number behave differently for independent and dependent degrees of freedom.
IV.1 Product states in fermion-boson hybrid descriptions
For defining entanglement of dependent degrees of freedom, we first need to recapitulate nonentangled states for independent degrees of freedom. Specifically, product states of independent fermions and bosons take the form
@@ IV.1 Product states in fermion-boson hybrid descriptions
where $\tilde{x}_{m}=x_{m}$ and $\tilde{y}_{n}=y_{n}/\sqrt{n!}$ relate expansion coefficients in the Fock basis and excitations of quantized fields. As examples, the basis states in Eq. (11) are classical states with respect to the fermion-boson separation. A state $|\psi\rangle$ that cannot be expressed as defined in Eq. (15) is, by definition, entangled.
Note that, throughout most parts of this work, we focus on pure states. However, it is worth mentioning that classically correlated states, i.e., separable states [34], can be introduced by using an ensemble of pure product states in the form of a density matrix. Still, all concepts discussed here can be extended to mixed, inseparable states, as discussed in Sec. II.
IV.2 Field-product states for dependent degrees of freedom
Generalizing the second line in Eq. (15) renders it possible to define nonentangled states for the quasiparticle modes $\hat{p}_{\pm}$, Eq. (8). That is, a state is said to be nonentangled in this $\pm$ description if it can be expressed in a factorized manner as follows:
@@ IV.2 Field-product states for dependent degrees of freedom
with complex coefficients $x_{m}$ and $y_{n}$ such that the state is normalized, ${}_{\pm}\langle x,y|x,y\rangle_{\pm}=1$. In the next section, we show that this operational definition exactly coincides with the one for dependent degrees from Sec. II.
According to the definition in Eq. (16), the following $N$-particle states, where $N=m+n$, are product states:
@@
Note that such states are generally product states with respect to the quasiparticle modes $\hat{p}_{\pm}$ but entangled with respect to the fermion-boson separation $\mathrm{FB}$ since a global mode transformation is used in Eq. (8). Also, we can choose $m,n\in\mathbb{N}$ since neither $\hat{p}_{+}^{{\dagger}m}$ nor $\hat{p}_{-}^{{\dagger}n}$ becomes zero for a finite integer, contrasting pure fermions.
IV.3 Linear dependence and disentangled NOON states
It is straightforward to observe that, for any two pairs of integers $(m,n)$ and $(m^{\prime},n^{\prime})$, which obey $(m\varphi^{2}-n\beta^{2})/\sqrt{m+n}\neq(m^{\prime}\varphi^{2}-n^{\prime}\beta^{2})/\sqrt{m^{\prime}+n^{\prime}}$, the states $|m,n\rangle_{\pm}$ and $|m^{\prime},n^{\prime}\rangle_{\pm}$ are linearly independent. However, for a fixed $N>0$, there are only two linearly independent elements in total since the $\mathrm{FB}$ expansion is based on two basis elements only [see Eq. (18)]. Therefore, the identities
@@ IV.3 Linear dependence and disentangled NOON states
Figure 2: NOON-type states from Eq. (22) on the circle in an FB coordinate system, for $N=5$ quasiparticles, where $\varphi=\beta=1/\sqrt{2}$ and $\vartheta=0$. All nonentangled states in the $\pm$ description, Eq. (21), are shown as bullet points, including a $\pm$ coordinate system of the dependent degrees of freedom, which is indicated through dashed lines. Note that $N+1=6$ states, plus their reflections about the origin, are shown in the real plane, ignoring complex phases of $\psi_{0}$ and $\psi_{N}$ for simplicity. For independent degrees of freedom, all but the states $|N,0\rangle_{\pm}$ and $|0,N\rangle_{\pm}$ (and their reflections $-|N,0\rangle_{\pm}$ and $-|0,N\rangle_{\pm}$) alone would be entangled.
@@
with $\psi_{0}\psi_{N}\neq 0$, are not necessarily entangled for $N>2$. Recall that, by contrast, all NOON states are entangled for independent degrees of freedom. Here, whenever the coefficients $\psi_{0}$ and $\psi_{N}$ take the form of Eq. (21), such NOON states clearly admit a product form $|m,n\rangle_{\pm}$ for the dependent degrees of freedom $\pm$. Again, we emphasize that this disentanglement of quasiparticle NOON states is distinct from common discussions of entanglement (see Fig. 2).
IV.4 Quantum-correlated eigenstates of noninteracting Hamiltonians
Next, we consider the noninteracting case, where $\kappa=0$ for the Hamiltonian in Eq. (6), and the independent degrees of freedom $\mathrm{FB}$. This yields a Hamiltonian which decomposes into the single-mode parts $\hbar\omega_{F}\hat{f}^{\dagger}\hat{f}$ and $\hbar\omega_{B}\hat{b}^{\dagger}\hat{b}$, as well as into vanishing interactions, $\hat{H}_{\mathrm{int}}=0$. In that case, the eigenstates are necessarily products, $|m,n\rangle_{\mathrm{FB}}$ [Eq. (11)], for the nondegenerate case $\delta\neq 0$. This is generally true for independent degrees of freedom and $\hat{H}=\hat{H}_{1}\otimes\hat{1}_{2}+\hat{1}_{1}\otimes\hat{H}_{2}$, with subsystem identities $\hat{1}_{j}$ and local Hamiltonians $\hat{H}_{j}$ for $j\in\{1,2\}$. Further, the quasiparticle modes $\hat{p}_{\pm}$ were introduced to achieve a noninteracting representation of the Hamiltonian for $\kappa\neq 0$ [Eq. (10)]. So the expectation derived from independent degrees of freedom would be that the $N$-particle eigenstates in Eq. (12) are products with respect to the $\pm$ separation.